# Geometric properties of the additional third-order transitions in the two-dimensional Potts model

### v2

@@ 4 Conclusion
This study establishes the existence of third-order phase transitions in the Potts model from a geometric perspective and identifies the positions of both the third-order dependent and independent transitions using two key order parameters: isolated spins and the average perimeter. Comparisons between the microcanonical inflection points and transition positions determined by specific heat under canonical conditions show strong agreement. Additionally, by analyzing the relationship between the largest Cluster Size (LCS) and the magnetization $\langle M\rangle$, the order of phase transitions was identified for different values of $q$. The results show that for $q=3$ and $q=4$, the system undergoes continuous (second-order) transitions, accompanied by a third-order dependent transition. However, when $q=6$ and $q=8$, the system exhibits discontinuous (first-order) transitions, where the third-order dependent transition vanishes. Regardless of the value of $q$, a third-order independent transition persists, serving as a precursor to the impending disruption of the system’s highly ordered state. This comprehensive investigation validates the robustness of these geometric order parameters in detecting higher-order transitions in complex systems, further enhancing our understanding of the intricate behavior of the Potts model across varying conditions.
5 Acknowledgments
@@
We would like to thank Professor Ying Tang for his valuable discussion. This work is supported by the National Natural Science Foundation of China (Grant No.12304257).

### v3

@@ IV Conclusion
This study establishes the existence of third-order transitions in the Potts model from a geometric perspective and identifies the positions of both the third-order independent and dependent transitions using two order parameters: isolated spins number and the average perimeter. Comparisons between the microcanonical inflection points and transition positions determined by specific heat under canonical conditions show strong agreement. Furthermore, by analyzing the relationship between the Largest Cluster Size (LCS) and the magnetization $\langle M\rangle$, the order of phase transitions was identified for different values of $q$. The results show that for $q=3$ and $q=4$, the system undergoes continuous (second-order) transitions, whereas for $q=6$ and $q=8$, it exhibits discontinuous (first-order) transitions.
The number of isolated spins in the ordered phase usually reaches a peak, regardless of the value of $q$. This indicates that, independent of the order of the system’s main phase transition, a third-order independent transition usually exists, serving as a precursor to the impending disruption of the highly ordered state of the system. In the disordered phase of the system, we observed an inflection point in the rate of change of the average cluster perimeter, which corresponds to the existence of a third-order dependent transition. However, when the system undergoes a first-order phase transition, the dramatic changes in the clusters result in the absence of any higher-order phase transition signal in the system’s average cluster perimeter in the disordered phase.
Based on the current research findings, we can confirm that the third-order independent transition usually exists, regardless of the order of the system’s main phase transition. However, for the third-order dependent transition, when the phase transition from the ordered phase to the disordered phase is second-order, a third-order dependent transition is present in the system. But when the phase transition from the ordered phase to the disordered phase is first-order, the signal of the third-order dependent transition vanishes. Therefore, when the system undergoes a first-order main phase transition, whether a higher-order dependent transition exists in the disordered phase remains to be further studied and verified.
V Acknowledgments
@@
I would like to express my sincere gratitude to Professor Ying Tang for the invaluable discussions that significantly enriched this work and provided profound insights throughout the research process. This work is supported by the National Natural Science Foundation of China (Grant No.12304257).